OpenStudy (anonymous):

HELPPPPP(: 6,9,12,15,18 3,7,11,15,19 10,15,20,25,30 FOR EACH ONE : Write a pattern rule Write an expression for the nth term (using n ex. 3n+1) use the expression to find the 60th term

OpenStudy (anonymous):

6+3*0,6+3*1,6+3*2,6+3*3.....6+3*(n-1) or 3(2+n-1) or 3(1+n)

OpenStudy (anonymous):

60th term for first sequence is 3(1+60)=180+3=183